Solve if possible 3x+y+2= 0 2 x-y+2=0 2x+y+2=0
Answer by ewatrrr(24785)


Hi,
Determining solution for this System of equations by graphing
3x+y+2= 0 OR y = -3x-2 (0,-2), (-2/3,0) ,(-1,1)
2x-y+2=0 OR y = 2x+2
2x+y+2=0 OR y = -2x-2
As alll 3 three lines will have no point in common:
There is NO solution for this system